# Primordial Black Holes in the Solar System

### v1

@@ 4.4 Extrapolating our results
Our simulations help us to investigate perturbations induced by PBHs for a certain range of masses $M_{\rm PBH}$ and timescales $t$. We find that the dependence on the mass of the PBH is only logarithmic as long as $b_{\rm max}\gg r$ and, for the parameters that we studied, the results are well described by the impulse model, discussed in Section 4.1. Therefore one can attempt to use our analytical model to extend the study to longer timescales and to larger PBH masses that we could not investigate so far due to the computational expense of running a large number of additional simulations. However, it is necessary to make one modification to the impulse model. The reason is that for large PBH masses and small timescales one can no longer take the integration boundaries in Equation 5 to infinity as the timescale of a fly-by of a black hole will take longer than the the observation time $t$. If one takes into account the finite distance travelled by the PBH one obtains:
@@ 4.5 Prospects for detecting PBHs
The quantity of interest for detecting PBHs is the ratio between the perturbations $|\delta\mathbb{r}|$ that they induce and the observational residual $\sigma_{r}$ in the distance between Earth and a given SSO. In Section 4.2 we found that the distribution of perturbations $p(|\delta\mathbb{r}|/r)$ is highly skewed. This means that the mean perturbation strength is not a good measure for detectability as the likelihood to observe it can be low. Of the 500 simulation that were performed for $t=20$ yr, 113 end up with a perturbation strength $|\delta\mathbb{r}|/r$ which is larger than the mean $\langle|\delta\mathbb{r}|/r\rangle$, a fraction of $\sim 27\%$. Therefore it is more useful to compare the median perturbation strength to the observartional residual. If $|\delta\mathbb{r}|_{\rm median}>\alpha\sigma_{r}$ then there is 50% chance for an observation with a signal-to-noise ration of $\alpha$. In Section 4.1 we have derived the impulse model to describe the mean perturbation strength. To a good approximation, the median perturbation strength can be estimated by setting the minimum impact parameter to $b_{\rm min}$ (Equation 10) when performing the integration over $b$,
@@
The reason is that the median perturbation will be dominated by the smallest ’typical’ impact parameter, whereas the mean takes into account smaller values of $b$ that can occur rarely and is thus skewed to larger values. We demonstrate the validity of this assumption in the left panel of Figure 3 where we show the result from Equation 15 as the green line, which agrees well with the median from the simulation data (black, dashed line).
@@
As a conservative estimate, we regard the perturbations induced by the halo of PBHs as a random signal and thus a SNR greater than 1 is required for detection. This is different from the approach by Tran et al. (2023) who investigate individual encounters where the signal can be predicted and thus extracted from the data for SNRs much smaller than 1, e.g. by template matching techniques. We find that a significant detection ($|\delta\mathbb{r}|_{\rm median}/\sigma_{r}\geq 3$) cannot be reached with 100 years of observational data. If the uncertainty in the distance $\sigma_{r}$ decreases by a factor of 30, then only around 20-30 years of data is required to detect PBHs with a mass $M_{\rm PBH}\in[10^{19},10^{24}]$ g with $3\sigma$ confidence (this corresponds to the 0.1 contour in Figure 4). Note that $|\delta\mathbb{r}|\sim\rho_{\rm CDM}$ and hence the detection window becomes larger if the local dark matter density is greater than the assumed value of $\rho_{\rm CDM}=7\times 10^{-25}$ g cm-3. We want to emphasize that the results presented in Figure 4 are based on the analytical impulse model and for particular parameters one should perform a set of numerical simulations to obtain more accurate results. It would also be interesting to study the probability of detection for a given SNR as a function of $M_{\rm PBH}$ and $t$. However, this requires a better understanding of $p(|\delta\mathbb{r}|/r)$ that is beyond the scope of this work.
@@ 4.6 Limitations of this work
Let us first emphasize that our work focuses on the cumulative gravitational effect that a halo of PBHs has on the solar system and does not consider the detectability of individual encounters. The latter has been studied by Tran et al. (2023) who find a strong mass dependence for the expected event rate.
With our model one is able to make general statements about the strength and evolution of perturbations, induced by PBHs and to study whether they are in conflict with current data. However, in order to detect individual PBHs, one needs a precise model for solar system ephemerides that includes more solar system bodies and treatment of finite-size effects, radiation pressure, relativistic corrections and other effects that we have neglected. A combination of such a model, together with a sophisticated analysis of solar system ephemerides could be a viable pathway to the detection of asteroid-mass PBHs.
For the halo of PBHs we have assumed a monochromatic mass function for simplicity. In reality, PBHs are likely to form with an extended mass distribution. While one can easily expand our analysis for a distribution of PBH masses, it is not straightforward to connect this distribution to the initial PBH mass function at formation in the early Universe. The main reason for this is that one has to deal with migration effects, where the heavier black holes sink towards the center of the halo while the lighter ones move outwards. In general, we do not expect that an extended mass function would change the perturbations of PBHs in a drastic way as we found little mass dependence in our results.
A more sizeable impact on our results could be expected if PBHs are clustered (see Section II of Carr et al. (2024) and references therein). If the size of such a cluster is small compared to the impact parameter, then it will simply act as a much heavier black hole. However, if the size of such a cluster is comparable to the solar system then our results could change in a nontrivial way. In addition, the degree of clustering might be dependent on the black hole mass in which case an extended mass function could have a sizeable impact on the result. Therefore, it would be worthwhile to study this aspect in a future work.
@@ 5 Summary
Our main finding is, that the perturbations induced by a halo of PBHs on the solar system are too small to be detectable with the current solar system data. We obtain this result by performing numerical simulations of the solar system, embedded in a DM halo of PBHs. The simulation results are used to quantify the perturbations, induced by the PBHs on the distances between the Earth and other planets. First, it is demonstrated that the strength of the perturbations depends only weakly on the PBH mass in the range $10^{18}\,\rm{g}<M_{\rm PBH}<10^{21}\,\rm{g}$. The physical reason for this is the fact that the strength of the gravitational force ($\sim M_{\rm PBH}$) is balanced by the scattering rate ($\sim 1/M_{\rm PBH}$). We then show for $M_{\rm PBH}=10^{20}$ g that after a time span of 20 years the perturbations are still more than an order of magnitude below the current precision of observational data for the Earth-Mars distance, which has been available for a similar time period. We also demonstrate that the simple idea of approximating the gravitational influence of PBHs on the solar system by variations of the central solar mass fails and predicts perturbation strengths that are orders of magnitudes too large.
Our results can be interpreted by a simple analytical model that provides an independent test of the accuracy of our simulation. In addition, it enables us to extrapolate to longer timescales and to a larger range of PBH masses. We find that the precision of solar system ephemerides has to increase by more than an order of magnitude for the effect of the PBH halo to become noticeable within a decade of observation. Our results challenge recent work that has excluded PBHs as a dark matter candidate, based on solar system ephemerides.
@@ Appendix A Relation between the perturbation strength and the impact parameter
In Section 4.1 we derived expressions for the relation between the perturbation strength $|\delta\mathbb{r}|/r$ and the impact parameter $b$ in the limit $b\gg r$ (Equation 6) and for $b\ll r$ (Equation 7). In the intermediate case $b\sim r$ the result will strongly depend on the precise geometry of the encounter. To investigate this effect we uniformly sample $10^{7}$ PBH positions in a plane [^1] We assume that the PBHs move perpendicular to the ecliptic plane around two SSOs that are separated by distance $r=1$. For each PBH position we determine the minimum of the distance to both bodies $b$ and the perturbation strength $|\delta\mathbb{r}|/r$. Figure 5 shows the range of values obtained for $|\delta\mathbb{r}|$ for each value of $b$ as the shaded area and the mean values as points. We have set $12GM_{\rm PBH}\Delta t/v_{\rm PBH}=r=$ 1 as we only care about the functional dependence on $b$. The analytical relation (Equation 8) that we use in this work is shown as the orange line and agrees well with the numerical data. The strongest deviation is a factor of 2.2 for $b\approx r/2$. If we integrate over the numerical relation then we find a deviation of at most 20%. This justifies using Equation 8 to describe the simulation results.

### v2

@@ 4.3 Extrapolating our results
Our simulations help us to investigate perturbations induced by PBHs for a certain range of masses $M_{\rm PBH}$ and timescales $t$. We find that the dependence on the mass of the PBH is only logarithmic as long as $b_{\rm max}\gg r$ and, for the parameters that we studied, the results are well described by the impulse model, discussed in Section 4.1. Therefore one can attempt to use our analytical model to extend the study to longer timescales and to wider range of PBH masses that we could not investigate so far due to the computational expense of running a large number of additional simulations. However, it is necessary to make one modification to the impulse model. The reason is that for large PBH masses and small timescales one can no longer take the integration boundaries in Equation 4 to infinity as the timescale of a fly-by of a black hole will take longer than the the observation time $t$. If one takes into account the finite distance travelled by the PBH one obtains:
@@ 4.4 Prospects for detecting PBHs
The quantity of interest for detecting PBHs is the ratio between the perturbations $|\delta\mathbb{r}|$ that they induce and the observational residual $\sigma_{r}$ in the distance between Earth and a given SSO. In Section 4.2 we found that the distribution of perturbations $p(|\delta\mathbb{r}|/r)$ is highly skewed. This means that the mean perturbation strength is not a good measure for detectability as the likelihood to observe it can be low. Of the 500 simulation that were performed for $t=20$ yr, 113 end up with a perturbation strength $|\delta\mathbb{r}|/r$ which is larger than the mean $\langle|\delta\mathbb{r}|/r\rangle$, a fraction of $\sim 27\%$. Therefore it is more useful to compare the median perturbation strength to the observational residual. If $|\delta\mathbb{r}|_{\rm median}>\alpha\sigma_{r}$ then there is 50% chance for an observation with a signal-to-noise ration of $\alpha$. In Section 4.1 we have derived the impulse model to describe the mean perturbation strength. To a good approximation, the median perturbation strength can be estimated by setting the minimum impact parameter to $b_{\rm min}$ (Equation 9) when performing the integration over $b$,
@@
The reason is that the median perturbation will be dominated by the smallest ’typical’ impact parameter $b_{\rm min}$, whereas the mean takes into account smaller values of $b$ that can occur rarely and is thus skewed to larger values. We demonstrate the validity of this assumption in the left panel of Figure 3 where we show the result from Equation 14 as the green line, which agrees well with the median from the simulation data (black, dashed line).
@@
As a conservative estimate, we regard the perturbations induced by the halo of PBHs as a random signal and thus a SNR greater than 1 is required for detection. We find that a significant detection ($|\delta\mathbb{r}|_{\rm median}/\sigma_{r}\geq 3$) cannot be reached with 100 years of observational data. If the uncertainty in the distance $\sigma_{r}$ decreases by a factor of 30, then only around 20-30 years of data are required to detect PBHs with a mass $M_{\rm PBH}\in[10^{19},10^{24}]$ g with $3\sigma$ confidence (this corresponds to the 0.1 contour in Figure 4). Note that $|\delta\mathbb{r}|\sim\rho_{\rm CDM}$ and hence the detection window becomes larger if the local dark matter density is greater than the assumed value of $\rho_{\rm CDM}=7\times 10^{-25}$ g cm-3. We want to emphasize that the results presented in Figure 4 are based on the analytical impulse model and for particular parameters one should perform a set of numerical simulations to obtain more accurate results. It would also be interesting to study the probability of detection for a given SNR as a function of $M_{\rm PBH}$ and $t$. However, this requires a better understanding of $p(|\delta\mathbb{r}|/r)$ that is beyond the scope of this work.
So far, we have treated the perturbations as a random signal and thus required a SNR$>$1 for detection. However, the perturbations induced by the encounters of PBHs likely have certain characteristics that can be exploited to make predictions of the expected signal $\delta r(t)$. In that case, one can attempt to use template matching techniques in order to detect perturbations with an amplitude well below the noise level $\sigma_{r}$. This was studied by Tran et al. (2023) who found that there will be a sizeable number of detectable events if the extraction of signals with a SNR well below $10^{-2}$ is achieved. A possible caveat regarding the detection of PBHs are potential degeneracies between the perturbations induced by the PBHs and other gravitational effects in the solar system. Accurate models of solar system ephemerides are obtained by iterating over a large number of free parameters, including the physical parameters of all SSOs. To assess the prospects of detecting PBHs one has to study the degree of degeneracy between the perturbations that they induce and a change in one of the other parameters of these models.
4.5 Poissonian Fluctuations
In our work, we analyze at the cumulative gravitational effect of PBHs on the orbits of solar system objects. We find that the perturbations induced by the black holes are dominated by the closest encounter (see Section 4.1). This is in contrast to the work by Loeb (2024) which studies the influence of Poissonian fluctuations in the PBH density. If the dark matter is composed of asteroid-mass black holes then their number within the solar system can undergo sizeable Poisson fluctuations. Loeb (2024) considered the rate of change in the total mass of PBHs within a distance $R$ from the Sun, given by
In his work it is then assumed that such fluctuations would induce similar effects as if a point mass in the center of the solar system would change its mass with the same rate, effectively changing the solar mass by a small amount. For $R=50$ AU and $M_{\rm PBH}\in[6\times 10^{18},10^{22}]$ g, this would lead to a rate $\delta\dot{M}$ larger than the constraint on $\dot{M}_{\rm sol}$ by Pitjeva et al. (2021) and thus he concluded that PBHs in this mass range cannot make up the entire dark matter.
It is important to note that these Poissonian fluctuations can only occur if there is at least one PBH in the volume of interest. For Mars one requires $M_{\rm PBH}\lesssim 3\times 10^{16}$ g to have at least one PBH inside its orbit. At these low PBH masses the fluctuations that follow from Equation 15 are well below the observational constraint from Pitjeva et al. (2021) and thus cannot be used to rule out PBHs as dark matter. For the outer planets, the fluctuations from Equation 15 can exceed the constraint on $\dot{M}_{\rm sol}$, as e.g. for Neptune one requires $M\lesssim 3\times 10^{20}$ g which implies a value of $\delta\dot{M}<2.5\times 10^{-13}\,\rm{M}_{\odot}\,\rm{yr}^{-1}$. However, it is not justified to apply the constraint by Pitjeva et al. (2021) for the outer planets in isolation as it was derived from the entire set of solar system data, assuming no dependence of $\dot{M}_{\rm sol}$ on $r$, and is dominated by the effect of fluctuations on the inner planets. On the one hand, this is because the observational data of the inner planets currently has the highest accuracy. On the other hand, a change in the solar mass has a much stronger effect on the inner planets compared to the outer ones, as the additional force from the change in the solar mass mass will be proportional to $\delta\dot{M}/r^{2}$. In essence, an observational constraint on $\dot{M}_{\rm sol}$ at the orbital distance of Neptune would be orders of magnitude weaker than the limit which Loeb (2024) is using. Note that similar remarks were already made by Cline (2024).
For the reasons stated above, one cannot rule out PBHs based on the Poisson fluctuation in their number density. Nevertheless, it remains of interest to compare the effects of these fluctuations to those from individual encounters. To investigate this, we perform a simulation where, at each timestep and for each planet, we count the number of PBHs $N_{\rm PBH}(t,r_{i})$ within the orbital distance $r_{i}$ of the planet $i$ to the Sun. We then set the solar mass for each planet to $M_{\rm sol,i}(t)=\mathrm{M}_{\odot}+M_{\rm PBH}N_{\rm PBH}(t,r_{i})$. The PBHs do not exert any gravitational force on the solar system bodies in this simulation, acting only through the fluctuations of the solar mass. One has to have at least one PBH inside the orbit of a planet to observe any fluctuations. For Mercury this implies $M_{\rm PBH}\lesssim 4\times 10^{14}$ g which would lead to a computationally infeasible number of PBHs within the entire solar system. We therefore study the inner and outer planets in two separate simulations. For the simulation containing the inner planets and the Earth’s moon we choose $M_{\rm PBH}=3\times 10^{13}$ g and $L_{\rm box}=10$ AU, whereas we take $M_{\rm PBH}=10^{17}$ g and $L_{\rm box}=150$ AU for the outer planets. This means that there will be on average $\langle N_{\rm PBH}\rangle\in[13,120,347,858]$ PBHs inside the orbits of Mercury, Venus, Earth, and Mars. For Jupiter, Saturn, Uranus, and Neptune there will be $\langle N_{\rm PBH}\rangle\in[12,89,737,2622]$ PBHs inside their orbits. In Figure 5 we display the perturbations of the vector between the Sun and each planet. The left panel shows the results for the inner solar system at $t=3$ yr, whereas the simulation for the outer planets are shown in the right panel for $t=1$ yr. Blue points indicate the median value for simulations where we include the gravity of the PBHs, whereas the orange points show the results for the simulations with a fluctuating solar mass. The perturbations induced by PBHs on the orbit of the planets are roughly an order of magnitude larger compared to the effect of Poisson fluctuations of the solar mass. Based on these results alone, one can not distinguish whether the disagreement is because the Poisson fluctuations are a small contribution to the total effect of PBHs in the solar system or because one cannot treat the Poisson fluctuations of the PBH number density as a change in the solar mass. However, the analytical model introduced in Section 4.1 suggests that the perturbations induced by PBHs can be predicted with good accuracy by taking into account the closest encounters. We display the model for the median perturbation strength from Equation 14 as a dashed line in Figure 5 and it shows reasonable agreement for $b_{\rm max}=3b_{\rm min}$. This suggests that Poisson fluctuations likely play only a minor role in the overall gravitational effect of PBHs on the solar system bodies.
Figure 5: Perturbation of the vector between the Sun and the planets in the solar system. The left panel shows the results for a simulation of the inner solar system with $M_{\rm PBH}=3\times 10^{13}$ g at $t=3$ yr. In the right panel, we present the results for the outer solar system with $M_{\rm PBH}=10^{17}$ g at $t=1$ yr. Each data point corresponds to the median perturbation of 100 simulation runs. The errorbars indicate the range that includes 90% of the simulation results. The blue points correspond to our usual simulation setup. In orange, we show the results of simulations where the solar mass is changing according to the Poisson fluctuations of PBHs $\delta\dot{M}$, as explained in Section 4.5.
@@ 4.6 Limitations of this work
With our model one is able to make general statements about the strength and evolution of perturbations induced by PBHs and to study whether they are in conflict with current data. However, in order to detect individual PBHs one needs a precise model for solar system ephemerides that includes more solar system bodies and treatment of finite-size effects, radiation pressure, relativistic corrections, and other effects that we have neglected. It is crucial to study the possible degeneracy between the mentioned effects and the perturbations induced by the PBHs. A combination of such a model, together with a sophisticated analysis of solar system ephemerides could be a viable pathway to the detection of asteroid-mass PBHs.
For the halo of PBHs, we have assumed a monochromatic mass function for simplicity. In reality, PBHs are likely to form with an extended mass distribution. While one can easily expand our analysis for a distribution of PBH masses, it is not straightforward to connect this distribution to the initial PBH mass function at formation in the early Universe. The main reason for this is that one has to deal with migration effects, where the heavier black holes sink towards the center of the halo while the lighter ones move outwards. In general, we do not expect that an extended mass function would change the perturbations of PBHs in a drastic way as we found little mass dependence in our results.
A more sizeable impact on our results could be expected if PBHs are clustered (see Section II of Carr et al. (2024) and references therein). If the size of such a cluster is small compared to the impact parameter, then it will simply act as a much heavier black hole. However, if the size of such a cluster is comparable to the solar system then our results could change in a non-trivial way. In addition, the degree of clustering might be dependent on the black hole mass in which case an extended mass function could have a sizeable impact on the result. Therefore, it would be worthwhile to study this aspect in a future work.
@@ 5 Summary
In this work, we perform numerical simulations of the solar system, embedded in a DM halo of PBHs. The simulation results are used to quantify the perturbations induced by the PBHs on the distances between the Earth and other planets. First, it is demonstrated that the strength of the perturbations depends only weakly on the PBH mass in the range $10^{18}\,\rm{g}<M_{\rm PBH}<10^{21}\,\rm{g}$. The physical reason for this is the fact that the strength of the gravitational force ($\sim M_{\rm PBH}$) is balanced by the scattering rate ($\sim 1/M_{\rm PBH}$). We then show for $M_{\rm PBH}=10^{20}$ g that after a time span of 20 years the perturbations are still more than an order of magnitude below the current precision of observational data for the Earth-Mars distance, which has been available for a similar time period. Therefore, PBHs cannot be directly constrained based on solar system ephemerides, challenging the results of recent work.
Our results are interpreted by a simple analytical model that provides an independent test of the accuracy of our simulations. In addition, it enables us to extrapolate to longer timescales and to a larger range of PBH masses. We find that the precision of solar system ephemerides has to increase by more than an order of magnitude for the effect of the PBH halo to become noticeable within a decade of observation. In order to reliably detect PBHs it is necessary to employ accurate models of solar system ephemerides that include finite-size effects, relativistic corrections and other effects that were neglected in this work.
@@ Appendix A Relation between the perturbation strength and the impact parameter
In Section 4.1 we derived expressions for the relation between the perturbation strength $|\delta\mathbb{r}|/r$ and the impact parameter $b$ in the limit $b\gg r$ (Equation 5) and for $b\ll r$ (Equation 6). In the intermediate case $b\sim r$ the result will strongly depend on the precise geometry of the encounter. To investigate this effect we uniformly sample $10^{7}$ PBH positions in a plane around two SSOs that are separated by distance $r=1$. For each PBH position we determine the minimum of the distance to both bodies $b$ and the perturbation strength $|\delta\mathbb{r}|/r$. Figure 6 shows the range of values obtained for $|\delta\mathbb{r}|$ for each value of $b$ as the shaded area and the mean values as points. We have set $12GM_{\rm PBH}\Delta t/v_{\rm PBH}=r=$ 1 as we only care about the functional dependence on $b$. The analytical relation (Equation 7) that we use in this work is shown as the orange line and agrees well with the numerical data. The strongest deviation is a factor of 2.2 for $b\approx r/2$. If we integrate over the numerical relation then we find a deviation of at most 20%. This justifies using Equation 7 to describe the simulation results.